OpenStudy (anonymous):

a median family income was 16,000 in the year 1980 and about 43,000 in 2000.. what is the slope of the line passing through the points (1980, 16,000) and (2000, 43,000) m=

OpenStudy (anonymous):

I just need help figuring these out I'm practicing for a final next week

OpenStudy (anonymous):

The numbers in your brackets are your actual coordinates on your graph. So the first value in the brackets in your x-coordinate and the second your y-coordinate. So what is m? m= (y2-y1)/(x2/x1)

OpenStudy (anonymous):

does that help you?

OpenStudy (anonymous):

yes it does thank you

OpenStudy (anonymous):

im horrible at math so please don't laugh but I got 27,000 and 20.. is there another step that I take?

OpenStudy (anonymous):

If you are just looking for the gradient of the line, than no there is no extra step. Is this what you did? m= (16000-43000) / (1980-2000) m=1350x

OpenStudy (anonymous):

how ever, if you want to find the equation of the line of that slope than tehre is one extra step.

OpenStudy (anonymous):

no I thought I was supposed to subtract 16000 from 43000 and subtract 1980 from 2000 lol I can see I'm way off

OpenStudy (anonymous):

You should take your coordinates as they are written, whatever the second pair of cordinates are than that will be x2 and y2. The way you phrased the question y2 is 16000

OpenStudy (anonymous):

ok I think I understand now, I'm going to work it out and possibly get back with you later with my answer just to see if i got it

OpenStudy (anonymous):

thanks for your help

OpenStudy (anonymous):

no problem.